# GenQ: Quantization in Low Data Regimes with Generative Synthetic Data

### v2

@@ Abstract
In the realm of deep neural network deployment, low-bit quantization presents a promising avenue for enhancing computational efficiency. However, it often hinges on the availability of training data to mitigate quantization errors, a significant challenge when data availability is scarce or restricted due to privacy or copyright concerns. Addressing this, we introduce GenQ, a novel approach employing an advanced Generative AI model to generate photorealistic, high-resolution synthetic data, overcoming the limitations of traditional methods that struggle to accurately mimic complex objects in extensive datasets like ImageNet. Our methodology is underscored by two robust filtering mechanisms designed to ensure the synthetic data closely aligns with the intrinsic characteristics of the actual training data. In case of limited data availability, the actual data is used to guide the synthetic data generation process, enhancing fidelity through the inversion of learnable token embeddings. Through rigorous experimentation, GenQ establishes new benchmarks in data-free and data-scarce quantization, significantly outperforming existing methods in accuracy and efficiency, thereby setting a new standard for quantization in low data regimes.
@@ 4 Methodology
Figure 2: The overall image synthesis and filtering procedure of GenQ. \⃝raisebox{-0.9pt}{1} For DFQ synthesis, we directly use the label as the text prompt, and \⃝raisebox{-0.9pt}{2} we use several metrics to filter our out-of-distribution synthetic images. \⃝raisebox{-0.9pt}{3} For data-scarce synthesis, real images are used to optimize prompt $\{\mathtt{S}\}$ (Eq. (missing) 11). We then generate the synthetic images with the optimized prompt.
@@ 4.1 Data-Free GenQ
In our label prompt method, we generate the images relying solely on the label name of each class and the ability of Stable Diffusion (Fig. 2\⃝raisebox{-0.9pt}{1}). However, this operation does not leverage any prior knowledge in the original training data, which might cause a distribution shift from original training data to synthetic training data. To generate better-quality data for quantization, we propose model-dependent selection to leverage the prior knowledge embedded in the pre-trained full-precision model.
We propose a set of filtering mechanisms for model-dependent selection. We argue that the synthetic data selection can be treated as an Out-of-Distribution (OOD) detection problem [48]. However, contrary to OOD detection, we intend to select in-distribution data from the generated synthetic data for quantization purposes (Fig. 2\⃝raisebox{-0.9pt}{2}).
@@ 4.2 Data-Scarce GenQ
where $M$ is the number of all object classes. This method learns the token across multiple objects, aiming to characterize the whole dataset [^2] We provide more technical details in Appendix A. . Fig. 2\⃝raisebox{-0.9pt}{3} illustrates the data-scarce prompt optimization. After optimization, we directly used the learned token embedding and the class label to generate images and filter the output images using the same technique in Fig. 2\⃝raisebox{-0.9pt}{2}.
@@ 5.3 Data Transferability Evaluation
[Uncaptioned image] Figure 5: Accuracy vs # syn. data with data-free PTQ using Genie-M. Table 5: Cross-model evaluation on synthetic data. We compare the data transferability between (a) existing methods and (b) GenQ . (a) Genie-D & PSAQ-ViT Source model Res18 MBV2 Res18 64.86 48.37 MBV2 64.21 51.47 ViT-B-16 47.27 17.96 (b) GenQ Source model Res18 MBV2 Res18 65.72 54.32 MBV2 65.52 54.82 ViT-B-16 65.44 54.24

### v3

@@ Abstract
In the realm of deep neural network deployment, low-bit quantization presents a promising avenue for enhancing computational efficiency. However, it often hinges on the availability of training data to mitigate quantization errors, a significant challenge when data availability is scarce or restricted due to privacy or copyright concerns. Addressing this, we introduce GenQ, a novel approach employing an advanced Generative AI model to generate photorealistic, high-resolution synthetic data, overcoming the limitations of traditional methods that struggle to accurately mimic complex objects in extensive datasets like ImageNet. Our methodology is underscored by two robust filtering mechanisms designed to ensure the synthetic data closely aligns with the intrinsic characteristics of the actual training data. In case of limited data availability, the actual data is used to guide the synthetic data generation process, enhancing fidelity through the inversion of learnable token embeddings. Through rigorous experimentation, GenQ establishes new benchmarks in data-free and data-scarce quantization, significantly outperforming existing methods in accuracy and efficiency, thereby setting a new standard for quantization in low data regimes. Code is released at https://github.com/Intelligent-Computing-Lab-Yale/GenQ.
@@ 4 Methodology
[…]
@@ 4.1 Data-Free GenQ
[…]
@@ 4.2 Data-Scarce GenQ
[…]
@@ 5.2 Quantization Performance Evaluation
Figure 4: Evaluation of data-scarce GenQ .
Table 4: Evaluation of data-free PTQ on ViTs (top-1 accuracy (%))
@@ 5.3 Data Transferability Evaluation
Figure 5: Accuracy vs # syn. data with data-free PTQ using Genie-M.
Table 5: Cross-model evaluation on synthetic data. We compare the data transferability between (a) existing methods and (b) GenQ .
(a) Genie-D & PSAQ-ViT
Source model Res18 MBV2
Res18 64.86 48.37
MBV2 64.21 51.47
ViT-B-16 47.27 17.96
(b) GenQ
Source model Res18 MBV2
Res18 65.72 54.32
MBV2 65.52 54.82
ViT-B-16 65.44 54.24
@@ 6 Conclusion
Acknowledgment
This work was supported in part by CoCoSys, a JUMP2.0 center sponsored by DARPA and SRC, the National Science Foundation (CAREER Award, Grant #2312366, Grant #2318152), and the DoE MMICC center SEA-CROGS (Award #DE-SC0023198)
Appendix 0.A Technical Details
Textual Inversion in Stable Diffusion. In Sec. 4.2, we optimize a token embedding for data-scarce quantization. Specifically, this technique is based on the Textual Inversion [18] where a token embedding is optimized for accurate recreation of an object from 3-5 images. To optimize the token embedding, we freeze all other parts of the Stable Diffusion including all the parameters from the encoder, decoder, text encoder, and the denoising U-Net. The training objective is the same as the training of Stable Diffusion, i.e. diffuse the real images latent into noisy latent and learn how to denoise it except we only update the text embedding (cf. Eq. 11). Note that the original textual inversion is designed to find the token of the exact object while we have multiple images from multiple classes. Hence, our objective is to find some token embedding that can universally describe the traits of the ImageNet data, aiming to learn distinct distribution characteristics.
Implementation of Token Embedding Learning. To optimize $\mathtt{\{S\}}$, we optimize it for 50k iterations with a batch size of 32 and use Adam optimizer with a constant learning rate of $45e-$ 4, following the convention of implementation [18]. Note that for each batch, we construct 32 different prompts since each image may have a unique label name.
Appendix 0.B Prompt Template
We use the following template to generate the prompt:
  1. 1.
    photo of a {C}.
  2. 2.
    rendering of a {C}.
  3. 3.
    cropped photo of the {C}.
  4. 4.
    the photo of a {C}.
  5. 5.
    photo of a clean {C}.
  6. 6.
    photo of a dirty {C}.
  7. 7.
    dark photo of the {C}.
  8. 8.
    photo of my {C}.
  9. 9.
    photo of the cool {C}.
  10. 10.
    close-up photo of a {C}.
  11. 11.
    bright photo of the {C}.
  12. 12.
    cropped photo of a {C}.
  13. 13.
    photo of the {C}.
  14. 14.
    good photo of the {C}.
  15. 15.
    photo of one {C}.
  16. 16.
    close-up photo of the {C}.
  17. 17.
    rendition of the {C}.
  18. 18.
    photo of the clean {C}.
  19. 19.
    rendition of a {C}.
  20. 20.
    photo of a nice {C}.
  21. 21.
    good photo of a {C}.
  22. 22.
    photo of the nice {C}.
  23. 23.
    photo of the small {C}.
  24. 24.
    photo of the weird {C}.
  25. 25.
    photo of the large {C}.
  26. 26.
    photo of a cool {C}.
  27. 27.
    photo of a small {C}.